### v2

@@ 4 Properties of the proposed score
and since we know that the first term is strictly positive, it suffices to show that the expression inside the curly brackets is negative. We bring this to the following form:
[…]
hence it suffices to show that the numerator is negative, since the denominator is again strictly positive. This gives:
@@
We proceed similarly to derive an additional bound based on Expression (8). More precisely:
@@
so by noticing that the first term is strictly positive and the product should be positive, we require that the expression inside the curly brackets is negative. A bit of re-arrangement gives:
@@
thus we require that the numerator is negative, since the denominator is strictly positive. Hence, we have:
@@
[…]
@@
Figure 4: Plot of bounds of $p$ for $r=2$. The rectangular region $[6,8]\times[10,24]$ on the left subplot is zoomed in on the right, illustrating how the 2 boundary curves become parallel as $p$ increases.
The points of intersection of the boundary curves is (as expected) given by $p=2^{r+1}+1$ and solving one of Expressions (9) or (10) for the corresponding $k$ value. We can see from both Figures 3 & 4 that the two boundary curves become parallel as $k\rightarrow+\infty$. We can compute their gradient to validate this:
@@
Therefore, we can access the set of solutions (corresponding to the region where both bounds are satisfied) by considering the line $p=2k+\alpha$. The constant $\alpha$ can be computed by noticing that this line will pass through the midpoint of each line segment defined by the two boundary curves for large $k$. Hence, we need to compute the difference between the two boundary curves:
@@
$=\frac{(k+1)(k^{2}-1)^{r}-k^{2r+1}}{k^{r}(k-1)^{r}}+1\overset{k\rightarrow\infty}{\longrightarrow}2,$
since the first term in the numerator involves a $k$ term of degree $12r+$ 1 (which cancels with the negative term on the numerator) and a term $k^{2r}$ that simplifies to $1$ upon division. The rest of the terms tend to 0 as $k\rightarrow\infty$ so the only term that remains is the unit coefficient of $k^{2r}$, which means that for large $k$ the vertical distance between the two boundary curves is equal to two units. Their midpoint is just one unit away from each curve, thus we require:
@@
which is satisfied for $\alpha=r$. This is shown below using the binomial series expansion of $(1+1/k)^{r}$:
@@
[…]
The proof is completed by considering that $\operatorname*{\operatorname{MAXLEN}}$ is always at least equal to one; thus the maximum score for $p<2^{r+1}+1$ is equal to $p$, while for $p\geq 2^{r+1}+1$, the maximum score is equal to the expression that we mentioned at the beginning of the proof, evaluated at $k=k^{*}$. However, since $\operatorname*{\operatorname{MAXLEN}}$ can be less than $k^{*}$ and given that the maximum score expression is increasing for $k\in\mathbb{Z}^{+}_{\leq k^{*}}$, the maximum score of nominal outlyingness is attained for $k=\min\{\lfloor(p-r)/2\rfloor,\operatorname*{\operatorname{MAXLEN}}\}$. ∎
@@ 5.1 Simulation Study
In order to illustrate the validity of our method, we conduct a simulation study on artificial data. Our proposed framework, which we denote by SONO (Scores Of Nominal Outlyingness) is benchmarked against two frequent pattern mining algorithms, specifically Frequent Pattern Outlier Factor (FPOF) [21] and Frequent Pattern Isolation (FPI) [26]. These methods present certain similarities with SONO in the way they define outliers, yet they require specifying values for $\operatorname*{\operatorname{MAXLEN}}$ and for the minimum relative support threshold. The minimum relative support threshold is analogous to the expression $\sigma_{d}/n$ under our framework, where $n$ denotes the number of observations in the data.
We generate our artificial data set $\boldsymbol{X}$ using the following model:
@@
[…]
In the above, $\boldsymbol{X}_{j}$ denotes the $j$th variable, $p$ is the number of variables and $\ell$ is the number of categorical levels. The last $p-2$ variables are essentially sampled from a discrete uniform distribution, while the first level for the first two variables occurs with probabilities $q_{1}$ and $q_{2}$, respectively. We set $q_{1}=0.10$ and $q_{2}=0.05$ in order to have one level being significantly less frequent than the rest. We vary the values of $n$, $p$ and $\ell$ and allow these to take values $n=200,500,1000$, $p=3,5,7$ and $\ell=2,3,5$. This leads to 27 different scenarios, simulated a hundred times each (using a different seed number), thus yielding 2700 artificial data sets. For each data set, we run FPOF and FPI with a minimum relative support threshold of $1/\ell$ and $\operatorname*{\operatorname{MAXLEN}}$ determined by our method using the relevant functions from the R package fpmoutliers. SONO is run three times; once assuming equal probabilities across all levels for all variables, once with the true probabilities, and once with probability vectors $\left(q_{j}/2,(1-q_{j}/2)/(\ell-1),\ldots,(1-q_{j}/2)/(\ell-1)\right)^{\intercal}$, $j=1,2$, for variables $\boldsymbol{X}_{1},\boldsymbol{X}_{2}$ and equal probabilities across each level for the remaining $p-2$ variables. We record the average rank for the observations possessing the first level for any of $\boldsymbol{X}_{1}$ or $\boldsymbol{X}_{2}$, the estimated value of $\operatorname*{\operatorname{MAXLEN}}$ and the average runtime for each of the three algorithms across twenty-five repetitions. Since some observations may end up with identical scores, we use the highest rank to resolve any ties (a higher rank here refers to a larger score of outlyingness). We also subtract the scores obtained with FPOF from a unit for their ranks to be comparable to these computed using SONO and FPI (FPOF assigns higher scores to inliers). We finally set $r=1$ to be more consistent with FPOF and FPI and $\alpha=0.05$ when running SONO.
@@
We finally report the median runtime in seconds for our experiments, which heavily depends on the value $\operatorname*{\operatorname{MAXLEN}}$, in Table 2. The reported times are only for the case of probability levels being set equal across each variable to ensure a fair comparison between the three methods. As we can see, SONO is quicker than FPI and FPOF when $\operatorname*{\operatorname{MAXLEN}}=1$ and consistently faster than FPI for $\operatorname*{\operatorname{MAXLEN}}\geq 3$. The median runtime of SONO increases on a large scale as $\operatorname*{\operatorname{MAXLEN}}$ grows (potentially due to its current implementation not being fully optimised for speed), yet it does not take extremely long to produce results which are typically more reliable than these of FPOF and FPI, as seen previously. Finally, we notice that a $\operatorname*{\operatorname{MAXLEN}}$ value of seven was only attained in 181 out of the 900 data sets consisting of seven variables and only when the true underlying level probabilities were provided. This highlights the efficiency of using support-based pruning, leading to a substantially smaller runtime in approximately 80% of the data sets with seven variables.
@@
Table 1: Median runtime in seconds for SONO, FPI and FPOF for MAXLEN values of one up to six, assuming equal probabilities across variable levels. Smallest times per $\operatorname*{\operatorname{MAXLEN}}$ value are bolded.
5.2 Applications on real data
We implement our method and compare its performance to that of FPOF and FPI on four publicly available data sets from the UCI Machine Learning Repository. The data sets used are the Solar Flare, the Differentiated Thyroid Cancer Recurrence, the Lymphography and the Wisconsin Breast Cancer data sets. These data sets consist of mainly nominal variables and the few variables of a different type (if any) are removed in the cleaning process, so that our method can be applied. Notice that these data sets have previously been used for classification purposes [see for example 34, 15, 9, 5, 8]. The Lymphography and Wisconsin Breast Cancer data sets have also been used for anomaly detection purposes in multiple instances [for instance in 21, 29, 49, 2, 26, among others].
The number of observations for each data set upon cleaning, as well as the number of variables used for each data set are summarised in Table 2, together with the value of $\operatorname*{\operatorname{MAXLEN}}$ as determined by SONO. Notice that we set $\alpha=0.05$ and $r=1$ for SONO, while the probabilities were set to be equal across variable categories. Following the implementations of FPOF and FPI on the Lymphography and Wisconsin Breast Cancer data sets from the papers that introduced these algorithms, we set the minimum relative support threshold to be $0.1$. The value of $\operatorname*{\operatorname{MAXLEN}}$ was set to be the same as the one estimated by SONO for each data set. We also report the proportion of outliers as a percentage for each data set, assuming that instances from the minority class are outlying for the Thyroid Cancer, Breast Cancer and Lymphography data sets. For the Solar Flare data set, observations corresponding to moderate or severe solar flares are taken to be the outliers.
Data set Observations Variables $\operatorname*{\operatorname{MAXLEN}}$ Proportion of Outliers (%)
Solar Flare 1389 10 4 5.26%
Thyroid Cancer 383 13 1 28.20%
@@ 5.2 Applications on real data
Lymphography 148 18 1 4.05%
Table 2: Number of observations upon data cleaning, number of variables used for nominal outlier detection, $\operatorname*{\operatorname{MAXLEN}}$ values and proportion of outliers for real-world data sets used.
Following the recommendations of [12], we compare the performance of the three algorithms using three evaluation measures. The first one we use is the Detection Rate at Top $K\%$, defined as the proportion of outliers detected within the observations with the top $K\%$ of outlyingness scores. This is analogous to the Recall measure, so we denote this by $\mathcal{R}(K)$. Assuming $O$ is the set of outliers, $\mathcal{R}(K)$ is defined as follows:
@@
where $\text{rank}(o)$ denotes the rank of an outlying instance $o\in O$ in a sorted list of outlyingness scores (ranked in descending order, with a rank of one being the most outlying). We also compute the average rank for the outlying observations and Area Under the Receiver Operating Characteristic Curve (ROC AUC) value for each method. The ROC AUC value for a list of ranked scores is computed as follows:
@@
$\mathcal{S}(o,i)$ =$=\left\{\begin{array}[]{ll}1,&\text{if rank}(o)>\text{rank}(i)\\ \frac{1}{2},&\text{if rank}(o)=\text{rank}(i)\\ 0,&\text{if rank}(o)<\text{rank}(i)\\ \end{array}\right.,$
where $I$ is the set of inliers [20]. A value closer to one indicates perfect performance, as this implies that most outliers are ranked higher than most of the inliers. In contrast, a value closer to zero will occur when an algorithm assigns higher outlyingness scores to inliers. We report $\mathcal{R}(K)$, the average rank and the ROC AUC for the $K\%$ top-ranked observations for each data set and each algorithm in Table 3 and Figures 6 & 7 for several values of $K$.
@@
Based on the average rank of outliers in Table 3, we see that SONO outperforms FPOF and FPI on the Wisconsin Breast Cancer data, while it achieves better performance than FPOF on the Lymphography data set. Despite not being able to do as well as its competitors on the Solar Flare and the Thyroid Cancer data sets, it has still managed to recover a good amount of outliers, as can be seen in Figure 6. In fact, for the Solar Flare data set, we see that the Detection Rate $\mathcal{R}(K)$ is comparable and even higher than that of FPOF or FPI for $K\leq 30\%$. Similarly for the Thyroid Cancer Data set, the largest scores of outlyingness obtained with each of the three methods correspond in their majority to outliers. Moreover, the ROC AUC for SONO takes very similar values to the ones obtained with FPI and FPOF, as seen in Figure 7. Hence, in a practical application, where the proportion of outliers would be assumed to be less than $50\%$, SONO would be a competitive algorithm to the ones presented, being able to correctly detect over 70% of the outliers.
@@
Overall, SONO has exhibited comparable performance to FPI and FPOF. However, it is important to stress that SONO required far less user input than its competitors, since the value for $\operatorname*{\operatorname{MAXLEN}}$ was estimated from the data and none of the minimum support threshold values was input by the user. Potentially different values for the minimum relative support threshold could have produced much different results for FPI and FPOF, yet these are not explored here and in fact, no recommendations have been made regarding the selection of an optimal value for this hyperparameter. A final appealing feature of SONO is that we can interpret the scores obtained and see which variables have contributed to these using the additional concepts introduced in Section 3. We will not delve into the details of the itemsets that contributed in the scores presented for each data set but we will give some basic guidelines on how the contribution matrix $\boldsymbol{\mathrm{{C}}}$ and the Nominal Outlyingness Depth can facilitate their interpretation. For instance, the contribution matrices for the outliers of each data set are illustrated in Figure 8. As can be seen, the size of the sun spot area or whether there was any activity in that region in the last twenty-four hours seem to be the most influential variables for a solar flare to be flagged as possibly moderate or severe. Similarly, the stage, metastasis (M) classification and whether the patient has received radioactive iodine therapy are indicators of thyroid cancer recurrence, while marginal adhesion and mitoses can be used to detect cases of malignant breast tumor. Finally, the rate of lymphatic diminishment and the blockage of lymphatic sinuses can hint normal finds or cases of fibrosis in the Lymphography data set. We also plot the Nominal Outlyingness Depth against the scores for the outliers in the Solar Flare data set, for which $\operatorname*{\operatorname{MAXLEN}}$ was larger than one in Figure 9. Some observations have a zero depth and score, while most outliers have at least a depth of one and one observation even has a depth of three. The largest score observed for the Solar Flare data set corresponds to a depth of nearly two units, suggesting that contributions from itemsets of both unit and greater lengths have resulted in this substantial score, making this specific observation a probable outlier.
@@
Figure 9: Nominal Outlyingness Depth (NOD) against scores of nominal outlyingness for the outliers of the Solar Flare data set.

### v3

@@ 4 Properties of the proposed score
so it suffices to show that the expression inside the curly brackets is negative.
$\begin{aligned} \frac{1}{\left(k^{*}+1\right)\left(1+\frac{1}{k^{*}}\right)^{r}}-\frac{1}{p-k^{*}}=\\ =\frac{\left(p-k^{*}\right)-\left(k^{*}+1\right)\left(1+\frac{1}{k^{*}}\right)^{r}}{\left(k^{*}+1\right)\left(1+\frac{1}{k^{*}}\right)^{r}\left(p-k^{*}\right)}\\ \end{aligned},$
so we show the numerator is negative, since the denominator is strictly positive. This gives
@@
We proceed similarly to derive an additional bound based on Expression (8). More precisely
@@
so since the first term is strictly positive and the product must be positive, we require that the expression inside the curly brackets is negative. Re-arranging gives
@@
thus we require that the numerator is negative, since the denominator is strictly positive. Hence, we have
@@
Thus, $k^{*}$ should be an integer such that Expressions (9) & (10) are satisfied. We plot the above two bounds of $p$ that we obtained for varying $r$ in Figure 3 to get an idea of what the region of solutions looks like and include a zoomed in version for the case of $r=2$ just to get a better understanding of what is going on in Figure 4.
@@
Figure 4: Plot of bounds of $p$ for $r=2$. The rectangular region $6,8]\times[10,24]$ on the left subplot is zoomed in on the right, illustrating how the 2 boundary curves become parallel as $p$ increases.
The point of intersection of the boundary curves corresponds to $p=2^{r+1}+1$ and the two boundary curves become parallel as $k\rightarrow+\infty$. We validate this by computing gradients below
@@
Therefore, we can access the set of solutions by considering the line $p=2k+\alpha$ and evaluating the value of $\alpha$. The difference between the two boundary curves is
@@
$=\frac{(k+1)(k^{2}-1)^{r}-k^{2r+1}}{k^{r}(k-1)^{r}}+1\overset{k\rightarrow\infty}{\longrightarrow}2.$
Thus, for large $k$ the vertical distance between the two boundary curves is equal to two units. Their midpoint is just one unit away from each curve, thus we require
@@
which is satisfied for $\alpha=r$. This is shown below using the binomial series expansion of $(1+1/k)^{r}$
@@
giving $\alpha=r$, as required. Notice that using this gives k=(pr)/2k=(p-r)/2, which means that for non-integer or odd integer values of prp-r we do not get an integer value for $k$. It is easy to show that for these cases, the only integer $k^{*}$ such that the bounds for $p$ are satisfied is $k^{*}=\lfloor(p-r)/2\rfloor$, where $\lfloor\cdot\rfloor$ is the floor function ($\lfloor x\rfloor$ is the greatest integer $x^{\prime}$ such that $x^{\prime}\leq x$).
Finally, as $\operatorname*{L_{\text{max}}}\geq 1$, the maximum score for $p<2^{r+1}+1$ is equal to $p$, while for $p\geq 2^{r+1}+1$, the maximum score is equal to the expression that we mentioned at the beginning of the proof, evaluated at $k=k^{*}$. Since $\operatorname*{L_{\text{max}}}$ can be less than $k^{*}$ and given that the maximum score expression is increasing for $k\in\mathbb{Z}^{+}_{\leq k^{*}}$, the maximum score of nominal outlyingness is attained for $k=\min\{\lfloor(p-r)/2\rfloor,\operatorname*{L_{\text{max}}}\}$. ∎
@@ 5.1 Simulation Study
In order to illustrate the validity of our method, we conduct a simulation study on artificial data. Our proposed framework, SONO, is benchmarked against two frequent pattern mining algorithms, specifically Frequent Pattern Outlier Factor (FPOF) [22] and Frequent Pattern Isolation (FPI) [27]. These methods present certain similarities with SONO in the way they define outliers, yet they require specifying values for $\operatorname*{L_{\text{max}}}$ and for the minimum relative support threshold. The minimum relative support threshold is analogous to the expression $\sigma_{d}/n$ under our framework, where $n$ denotes the number of observations in the data.
We generate our artificial data set $\boldsymbol{X}$ using the model below
@@
$\boldsymbol{X}_{3},\ldots,\boldsymbol{X}_{p-1},\boldsymbol{X}_{p}\overset{\text{i.i.d.}}{\sim}\text{Categorical}\left(\underbrace{\frac{1}{\ell},\ldots,\frac{1}{\ell}}_{\ell\text{ times}}\right).$
In the above, $\boldsymbol{X}_{j}$ denotes the $j$th variable, $p$ is the number of variables and $\ell$ is the number of categorical levels. The last p2p-2 variables are essentially sampled from a discrete uniform distribution, while the first level for the first two variables occurs with probabilities $q_{1}$ and $q_{2}$, respectively. We set $q_{1}=0.10$ and $q_{2}=0.05$ in order to have one level being significantly less frequent than the rest. We vary the values of $n$, $p$ and $\ell$ and allow these to take values $n=200,500,1000$, $p=3,5,7$, and $\ell=2,3,5$. This leads to 27 different scenarios, simulated a hundred times each (using a different seed number), thus yielding 2700 artificial data sets.
For each data set, we run FPOF and FPI with a minimum relative support threshold of $1/\ell$ and $\operatorname*{L_{\text{max}}}$ determined by our method using the relevant functions from the R package fpmoutliers. SONO is run using the R package SONO [16] three times; once assuming equal probabilities across all levels for all variables, once with the true probabilities, and once with probability vectors $\left(q_{j}/2,(1-q_{j}/2)/(\ell-1),\ldots,(1-q_{j}/2)/(\ell-1)\right)^{\intercal}$, $j=1,2$, for variables $\boldsymbol{X}_{1},\boldsymbol{X}_{2}$ and equal probabilities across each level for the remaining p2p-2 variables. We record the average rank for the observations possessing the first level for any of $\boldsymbol{X}_{1}$ or $\boldsymbol{X}_{2}$, the estimated value of $\operatorname*{L_{\text{max}}}$ and the average runtime for each of the three algorithms across twenty-five repetitions. Since some observations may end up with identical scores, we use the highest rank to resolve any ties (a higher rank here refers to a larger score of outlyingness). We also subtract the scores obtained with FPOF from a unit for their ranks to be comparable to these computed using SONO and FPI (FPOF assigns higher scores to inliers). We finally set $r=1$ to be more consistent with FPOF and FPI and $\alpha=0.05$ when running SONO.
@@
We finally report the median runtime in seconds for our experiments, which heavily depends on the value $\operatorname*{L_{\text{max}}}$, in Table 2. The reported times are only for the case of probability levels being set equal across each variable to ensure a fair comparison between the three methods. As we can see, SONO is quicker than FPI and FPOF when $\operatorname*{L_{\text{max}}}=1$ and consistently faster than FPI for $\operatorname*{L_{\text{max}}}\leq 3$. The median runtime of SONO increases on a large scale as $\operatorname*{L_{\text{max}}}$ grows (potentially due to its current implementation not being fully optimised for speed), yet it does not take extremely long to produce results which are typically more reliable than these of FPOF and FPI, as seen previously. Finally, we notice that a $\operatorname*{L_{\text{max}}}$ value of seven is only attained in 181 out of the 900 data sets consisting of seven variables and only when the true underlying level probabilities are provided. This highlights the efficiency of using support-based pruning, leading to a substantially smaller runtime in approximately 80% of the data sets with seven variables.
@@
Table 1: Median runtime in seconds for SONO, FPI and FPOF for MAXLEN values of one up to six, assuming equal probabilities across variable levels. Smallest times per $\operatorname*{L_{\text{max}}}$ value are bolded.
5.2 Applications on publicly available data
We implement our method and compare its performance to that of FPOF and FPI on four publicly available data sets from the UCI Machine Learning Repository. The data sets used are the Solar Flare, the Differentiated Thyroid Cancer Recurrence, the Lymphography, and the Wisconsin Breast Cancer data sets. These data sets consist of mainly nominal variables and the few variables of a different type (if any) are removed in the cleaning process, so that our method can be applied. Notice that these data sets have previously been used for classification purposes [see for example 34, 15, 9, 5, 8]. The Lymphography and Wisconsin Breast Cancer data sets have also been used for anomaly detection purposes in multiple instances [for instance in 22, 30, 49, 2, 27, among others].
The number of observations for each data set upon cleaning, as well as the number of variables used for each data set are summarised in Table 2, together with the value of $\operatorname*{L_{\text{max}}}$ as determined by SONO. Notice that we set $\alpha=0.05$ and $r=1$ for SONO, while the probabilities are set to be equal across variable categories. Following the implementations of FPOF and FPI on the Lymphography and Wisconsin Breast Cancer data sets from the papers that introduced these algorithms, we set the minimum relative support threshold to be $0.1$. The value of $\operatorname*{L_{\text{max}}}$ is set to be the same as the one estimated by SONO for each data set. We also report the proportion of outliers as a percentage for each data set, assuming that instances from the minority class are outlying for the Thyroid Cancer, Breast Cancer and Lymphography data sets. For the Solar Flare data set, observations corresponding to moderate or severe solar flares are taken to be the outliers.
Data set Observations Variables $\operatorname*{L_{\text{max}}}$ Proportion of Outliers (%)
Solar Flare 1389 10 6 5.26%
Thyroid Cancer 383 13 4 28.20%
@@ 5.2 Applications on publicly available data
Lymphography 148 18 4 4.05%
Table 2: Number of observations upon data cleaning, number of variables used for nominal outlier detection, $\operatorname*{L_{\text{max}}}$ values and proportion of outliers for publicly available data sets used.
Following the recommendations of [12], we compare the performance of the three algorithms using three evaluation measures. The first one we use is the Detection Rate at Top $K\%$, defined as the proportion of outliers detected within the observations with the top $K\%$ of outlyingness scores. This is analogous to the Recall measure, so we denote this by $\mathcal{R}(K)$. Assuming $O$ is the set of outliers, $\mathcal{R}(K)$ is defined as follows
@@
where $\text{rank}(o)$ denotes the rank of an outlying instance $o\in O$ in a sorted list of outlyingness scores (ranked in descending order, with a rank of one being the most outlying).
We also compute the average rank for the outlying observations and Area Under the Receiver Operating Characteristic Curve (ROC AUC) value for each method. The ROC AUC value for a list of ranked scores is computed as follows
@@
$\mathcal{S}(o,i)$ $=\left\{\begin{array}[]{ll}1,&\text{if rank}(o)>\text{rank}(i)\\ \frac{1}{2},&\text{if rank}(o)=\text{rank}(i)\\ 0,&\text{if rank}(o)<\text{rank}(i)\\ \end{array}\right$.,
where $I$ is the set of inliers [21]. A value closer to one indicates perfect performance, as this implies that most outliers are ranked higher than most of the inliers. In contrast, a value closer to zero occurs when an algorithm assigns higher outlyingness scores to inliers. We report $\mathcal{R}(K)$, the average rank and the ROC AUC for the $K\%$ top-ranked observations for each data set and each algorithm in Table 3 and Figures 6 & 7 for several values of $K$.
@@
Based on the average rank of outliers in Table 3, we see that SONO outperforms FPOF and FPI on the Wisconsin Breast Cancer data, while it achieves better performance than FPOF on the Lymphography data set. The $\operatorname*{L_{\text{max}}}$ values for Thyroid Cancer, Wisconsin Breast Cancer, and Lymphography are also considerably smaller than the number of variables. This makes the search for infrequent itemsets less computationally intensive without necessarily affecting performance.
@@
Despite not being able to do as well as its competitors on the Solar Flare and the Thyroid Cancer data sets, SONO has managed to recover a good amount of outliers, as can be seen in Figure 6. In fact, for the Solar Flare data set, we see that the Detection Rate $\mathcal{R}(K)$ is comparable to that of FPOF or FPI for $K\leq 30\%$. We also observe an almost identical performance of the three methods on the Wisconsin Breast Cancer data set, whereas SONO outperforms its competitors on Lymphography.
The ROC AUC for SONO takes very similar values to the ones obtained with FPI and FPOF, as seen in Figure 7. A constant ROC AUC score of a unit is achieved by SONO on the Lymphography data set, indicating that all outliers in the data set received higher scores than the inliers. Finally, we highlight that in a practical application, where the proportion of outliers is typically assumed to be less than $50\%$, SONO is a competitive algorithm to the ones presented, being able to correctly detect over 70% of the outliers.
Overall, SONO exhibits comparable performance to FPI and FPOF. However, it is important to stress that SONO requires far less user input than its competitors, since the value for $\operatorname*{L_{\text{max}}}$ is estimated from the data and none of the minimum support threshold values are input by the user. Different values for the minimum relative support threshold may produce much different results for FPI and FPOF, yet these are not explored here and in fact, no recommendations are made regarding the selection of an optimal value for this hyperparameter.
@@
A final appealing feature of SONO is that we can interpret the scores obtained and see which variables have contributed to these using the additional concepts introduced in Section 3. For instance, the contribution matrices for the outliers of each data set are illustrated in Figure 8. As can be seen, the size of the sun spot area or whether there is any activity in that region in the last twenty-four hours seem to be the most influential variables for a solar flare to be flagged as possibly moderate or severe. Similarly, the stage, metastasis (M) classification, and whether the patient has received radioactive iodine therapy are indicators of thyroid cancer recurrence, while marginal adhesion and mitoses can be used to detect cases of malignant breast tumor. Finally, the rate of lymphatic diminishment and the blockage of lymphatic sinuses can hint normal finds or cases of fibrosis in the Lymphography data set.
We also plot the Nominal Outlyingness Depth against the scores for the outliers in the Solar Flare, Thyroid Cancer and Lymphography data sets, for which $\operatorname*{L_{\text{max}}}$ is larger than one in Figure 9. Some observations have a zero depth and score, while most outliers have at least a depth of one and one observation even has a depth of two. The estimated values of $\operatorname*{L_{\text{max}}}$ are much higher for all three data sets than the maximum nominal outlyingness depths attained. This hints that the longer itemsets do not contribute in the calculation of the scores; e.g. in the Solar Flare data set, where the maximum depth is equal to a unit. This highlights a drawback of our strategy for selecting $\operatorname*{L_{\text{max}}}$, as calculations for itemsets of length two up to six on Solar Flare are eventually redundant, at least for distinguishing the outliers from the rest of the observations.
@@
Figure 9: Nominal Outlyingness Depth (NOD) against scores of nominal outlyingness computed with SONO for the outliers of the Solar Flare, Thyroid Cancer and Lymphography data sets.